The mayor of a town has proposed a plan for the annexation of an adjoining bridge. A political study took a sample of 1700 voters in the town and found that 35% of the residents favored annexation. Using the data, a political strategist wants to test the claim that the percentage of residents who favor annexation is more than 32% . Make the decision to reject or fail to reject the null hypothesis at the 0.05 level.

A. Reject Null Hypothesis

B. Fail to Reject Null Hypothesis
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p = 0.32

Ha : p \neq 0.32

\hat p = x / n =

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.35 - 0.32 / [\sqrt(0.32 * 0.68) / 1700]

= 2.65

P-value = 0.004

\alpha = 0.05

P-value < \alpha

Reject the null hypothesis .

A